OpenStudy (anonymous):

what is the energy (Joules) of x-radiation with a wavelength of 1.00x10^-6m?

OpenStudy (anonymous):

v=frequency c=speed of light=3.00e8m/s wavelength = 1.00e-6m wave length= speed of light/frequency what you need to find is v=c/wavelength Once you find the frequency use: E=hv E=energy in joules h always equal 6.63e10-34J*s (Planck's Constant) v= the frequency you got from the first part